# Symmetry, Invariant Manifolds and Flow Reversals in Active Nematic Turbulence

### v1

@@ V.3 Chaotic Set $\mathcal{Y}$
V.3.1 Approximations to the ECSs of a Typical Trajectory in $\mathcal{Y}$
@@ Acknowledgments
Table 3: Isotropy groups of the unstable eigenspaces of UNI. All eigenvalues are complex. The $\mathrm{R}_{\mathrm{a}}$ column indicates the approximate activity values at which each two-dimensional unstable eigenspace emerges. From $\mathrm{R}_{\mathrm{a}}\approx 3.7$ to $4.5$, no additional local bifurcations from UNI are observed.
@@
3rd $\{\sigma_{x},\tau_{x}(L/2)\}$ $\{e,\sigma_{x},\tau_{x}(L/2),\sigma_{x}\,\tau_{x}(L/2)\}$ $D_{2}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 1.79
4th $\{\sigma_{x},\tau_{x}(L/3)\}$ $\{e,\tau_{x}(L/3),\tau_{x}(2L/3),\sigma_{x},\sigma_{x}\,\tau_{x}(L/3),\sigma_{x}\,\tau_{x}(2L/3)\}$ $\mathbb{Z}_{6}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{3}$ 2.03
5th $\{\sigma_{x}\,\tau_{x}(L/6)\}$ $\{e,\sigma_{x}\,\tau_{x}(L/6),\tau_{x}(L/3),\sigma_{x}\,\tau_{x}(L/2),$ $\tau_{x}(2L/3),\sigma_{x}\,\tau_{x}(5L/6)\}$ $\mathbb{Z}_{6}$ 2.5
@@
Table 4: Isotropy groups of the unstable eigenspaces of LAN. Each real eigenvalue has multiplicity two. The $\mathrm{R}_{\mathrm{a}}$ column indicates the approximate activity values at which each two-dimensional unstable eigenspace emerges.
@@
6th $\{\tau_{x}(L/2),\tau_{x}(L/4)\}$ $\{e,\tau_{x}(L/4),\tau_{x}(L/2),\tau_{x}(3L/4)\}$ $\mathbb{Z}_{4}$ 3.30
7th $\{\tau_{x}(L/2),\tau_{x}(L/4)\}$ $\{e,\tau_{x}(L/4),\tau_{x}(L/2),\tau_{x}(3L/4)\}$ $\mathbb{Z}_{4}$ 4.21
8th $\{\tau_{x}(L/2),\tau_{x}(L/4)\}$ $\{e,\tau_{x}(L/4),\tau_{x}(L/2),\tau_{x}(3L/4)\}$ $\mathbb{Z}_{4}$ 4.36

### v2

@@ V.3 Chaotic Set $\mathcal{Y}$
V.3.1 Dominant ECSs in $\mathcal{Y}$
@@ Acknowledgments
Table 3: Isotropy groups of the unstable eigenspaces of UNI in $L=L_{A}$. All eigenvalues are complex. The $\mathrm{R}_{\mathrm{a}}$ column indicates the approximate activity values at which each two-dimensional unstable eigenspace emerges. From $\mathrm{R}_{\mathrm{a}}\approx 3.7$ to $4.5$, no additional local bifurcations from UNI are observed.
@@
3rd $\{\sigma_{x}\tau_{x}(L/2),\tau_{x}(L/3)\}$ $\{e,\tau_{x}(L/3),\tau_{x}(2L/3),\sigma_{x}\,\tau_{x}(L/6),$ $\sigma_{x}\,\tau_{x}(L/2),\sigma_{x}\,\tau_{x}(5L/6)\}$ $D_{3}$ 1.79
4th $\{\sigma_{x},\tau_{x}(L/2)\}$ $\{e,\sigma_{x},\tau_{x}(L/2),\sigma_{x}\,\tau_{x}(L/2)\}$ $D_{2}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 2.05
5th $\{\sigma_{x},\tau_{x}(L/3)\}$ $\{e,\tau_{x}(L/3),\tau_{x}(2L/3),\sigma_{x},\sigma_{x}\,\tau_{x}(L/3),$ $\sigma_{x}\,\tau_{x}(2L/3)\}$ $\mathbb{Z}_{6}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{3}$ 2.5
@@
Table 4: Isotropy groups of the unstable eigenspaces of LAN in $L=L_{A}$. Each real eigenvalue has multiplicity two. The $\mathrm{R}_{\mathrm{a}}$ column indicates the approximate activity values at which each two-dimensional unstable eigenspace emerges.
@@
6th $\{\tau_{x}(L/4)\}$ $\{e,\tau_{x}(L/4),\tau_{x}(L/2),\tau_{x}(3L/4)\}$ $\mathbb{Z}_{4}$ 3.30
7th $\{\tau_{x}(L/4)\}$ $\{e,\tau_{x}(L/4),\tau_{x}(L/2),\tau_{x}(3L/4)\}$ $\mathbb{Z}_{4}$ 4.21
8th $\{e\}$ $\{e\}$ Trivial group 4.36